Review Part A A 1000 kg weather rocket is launched straight up. The rocket motor provides a constant acceleration for 16 s, then the motor stops. The rocket altitude 20 s after launch is 7100 m. You can ignore any effects of What was the rockets acceleration during the first 16 s? Express your answer with the appropriate units air resistance al x.10 | a= | Value Units Submit Previous Answers Request AnswerA student standing on the ground throws a ball straight up. The ball leaves the students hand with a speed of 17.0 m/s when the hand is 2.10 m above the ground How long is the ball in the air before it hits the ground? (The student moves her hand out of the way Express your answer with the appropriate units. You may want to review (Pages 49 - 51) For help with math skills, you may want to review: Quadratic Equations Value For general problem-solving tips and strategies for this topic, you may want to view a Video Tutor Solution of Time in the air for a tossed bal. Submit Prevnous Answers Request AnswerAs a science project, you drop a watermelon off the top of the Empire State Building, 320 m above the sidewalk. It so happens that Superman flies by at the instant you release the watermelon. Superman is headed straight down with a speed of 33.0 m/s How fast is the watermelon going when it passes Superman? Express your answer with the appropriate units Value Units
Answer #1

(1) Suppose acceleration of rocket for first 16 seconds = a m/s^2

So, speed of the rocket after 16 s is -

v = a*t = a*16 = 16a m/s

Distance traveled in 16 s -

s1 = (1/2)*a*t^2 = 0.5*a*16^2 = 128a meter.

Now, when the rocket engine shuts down after 16 s, gravitational acceleration 'g' will come in picture.

so -

s2 = v*t' - (1/2)*g*t^2

s2 = 16a*(20-16) - 0.5*9.8*(20-16)^2

=> s2 = 64a - 78.4 meter

Given that -

s1 + s2 = 7100 m

=> 128a + 64a - 78.4 = 7100

=> 192a = 7178.4

=> a = 7178.4 / 192 = 37.4 m/s^2 (Answer)

(2) Solution in enclosed sheet.

ちS 2-10 use eression - a.o 03S 9-8 veorti So, e expre sion - (2x+4) re value wer.